# Fast Flocking of Protesters on Street Networks

### v2

@@ 4.3 Best tactics
We now focus on the two main tactics that achieve flocking: the Alignment strict tactic and the best tactic (the one that corresponds to the rightmost dot on Figure 2). Figure 3 displays their scores over time in green and purple colors, respectively.
The plots show that the mobility score of both tactics first rapidly grow, and that this growth significantly decreases over time. Even if these tactics are not as good as the Propulsion strict tactic regarding mobility, they have comparable performances for this metric.
Notice that the Propulsion strict tactic has very low gathering scores (Figure 3, left), which makes it an irrelevant tactic despite its mobility score. Instead, both the Alignment strict tactic and the best static quickly reach excellent gathering scores. The best tactic significantly outperforms the Alignment strict tactic and has a linear gathering score plot in log-log scale. This means that its evolution has a polynomial growth (of exponent below $1$), indicating a fast growth, but also that the evolution of group size tends to flatten over time. This is due to groups reaching a state where all clusters of groups are in distant regions of the network. Then, it takes longer for groups to meet other groups, merge, and grow in size.
Finally, Figure 3 also displays the sprawling score for all considered tactics. We observe that the best tactic produces a greater sprawling score than Alignment. The sprawling, for those two tactics, is due to groups following each others when they detect another group on a neighbor node, without necessarily merging with it.
@@
As explained in Section 4.1, an efficient tactic should have a low sprawling score. The sprawling of the baseline is 1 (up to the third digit), thanks to the collective decision. This is the optimum value.
Our walkers do not have access to collective decision. When a cluster of groups arrives at an intersection, at the end of a street, it can split into multiple groups. In the case of a lone group (a cluster with no sprawling) it will split into multiple groups, with equal number of walkers on average. This reduces the gathering score.
@@ 4.4 Interpretation
Recall that the best tactic corresponds to the rightmost dot among those in the upper right corner of plots in Figure 2. It is defined by the following parameters: $\{\alpha_{follow}=0.1,\alpha_{align}=0.8,\alpha_{attr}=0.1\}$ and a null weight for other rules. This tactic produces groups of $121$ walkers on average, and walkers explore on average $540$ distinct nodes during their $1000$ steps. These scores are more than twice and five times more than what the baseline gets, respectively.
@@
Figure 4: Schematic configurations that walkers, groups and clusters of groups achieve with the best tactics.
Second, this same rule guarantees that, if two groups cross path, they then merge in a single cluster in which all walkers will follow the same path. Indeed when a group $u$ cross path with a smaller group $v$, we have $x_{i}(t+1)=x_{j}(t)$ and $x_{j}(t+1)=x_{i}(t)$ for all $i\in g_{u}(t)$ and all $j\in g_{v}(t)$. The net flux is then $\Delta J_{u\rightarrow v}(t)=n_{u}(t)-n_{v}(t)>0$ for all walkers, which drive them all in the same direction: the smaller group goes back towards $v$ (where it came from) while the trajectory of the largest group is left unchanged.
However, this rule alone is not perfectly effective at avoiding splitting. The walkers in the first group in a cluster will not always all choose the same node at an intersection, as illustrated in the two bottom pictures of Figure 4, as the different possible nodes all have a flux equal to 00. In this context, the Attraction rule allows walkers that split in the least chosen direction at an intersection to go backward and avoid loosing sight of the cluster.
@@ 5 Related work
Our approach is different from usual protest studies based on thresholds [7] or agent-based [8, 9, 10] model. Indeed, these works focus on how people decide to participate to a protest; they do not deal with protester mobility.
In distributed systems, computer scientists proposed solutions to gathering problems where walkers follow a common distributed algorithm to meet on any connected graph [11, 12].
Flocking is a collective dynamic where groups of walkers move spontaneously in the same direction [13, 14, 15]. It has been largely studied in free space. Few articles explore flocking when trajectories of walkers are network constrained. For example, in [16] the authors implement rules similar to ours to study the possible outcomes of their combinations on a line. They get results very similar to ours in a single street.
Still, to our knowledge, we are the first to experiment rules for flocking that we can apply on any network.

### v4

@@ 4.3 Best tactics
We now focus on the two main tactics that achieve flocking: the Alignment strict tactic and the best tactic (the one that corresponds to the rightmost dot on Figure 4). Figure 5 (left and center) displays their scores over time in green and purple colors, respectively.
The plots show that the coverage score of both tactics first rapidly grow, and that this growth significantly decreases over time. Even if these tactics are not as good as the Propulsion strict tactic regarding coverage, they have comparable performances for this metric.
The Propulsion strict tactic has very low gathering scores (Figure 5, left), which makes it an irrelevant tactic despite its coverage score. Instead, both the Alignment strict tactic and the best static quickly reach excellent gathering scores. The best tactic significantly outperforms the Alignment strict tactic and has a linear gathering score plot in log-log scale. This means that its evolution has a polynomial growth (of exponent below $1$), indicating a fast growth, but also that the evolution of group size tends to flatten over time. This is due to groups reaching a state where all clusters of groups are in distant regions of the network. Then, it takes longer for groups to meet other groups, merge, and grow in size.
Finally, Figure 5 (right) displays the sprawling score for all considered tactics. We observe that the best tactic produces a greater sprawling score than Alignment. The sprawling, for those two tactics, is due to groups following each others when they detect another group on a neighbor node, without necessarily merging with it.
@@
As explained when defining the metrics, an efficient tactic should have a low sprawling score. The sprawling of the baseline is 1 (up to the third digit), thanks to the collective decision. This is the optimum value.
Our walkers do not have access to collective decision, though. When a cluster of groups arrives at an intersection, at the end of a street, it may split into multiple groups. In the case of a lone group (a cluster with no sprawling) it will split into multiple groups, with equal number of walkers on average. This reduces the gathering score.
However, with a larger sprawling, a cluster keeps most of its groups unchanged after an intersection. Indeed, the first of its groups to reach the intersection node might split into multiple groups, as described above. However, if using an Alignment-based tactic, next walkers will then all follow the largest of the new groups that came from the split. This shows an effective tactic can benefit from a little sprawling to maintain its clusters walker-size.
@@ 4.4 Interpretation
Recall that the best tactic corresponds to the rightmost dot among those in the upper right corner of plots in Figure 4. It corresponds to the following parameters: $\{\alpha_{follow}=0.1,\alpha_{align}=0.8,\alpha_{attr}=0.1\}$ and a null weight for other rules. This tactic produces groups of $121$ walkers on average, and walkers explore on average $540$ distinct nodes during their $1000$ steps. These scores are more than twice and five times more than what the baseline gets, respectively.
(a) 1. Pushes forward
(b) 2. Merges clusters
(c) 3. At intersection…
(d) … little splitting
Figure 6: Typical behaviors of walkers, groups and clusters of groups with the best tactic.
@@
Second, this same rule guarantees that, if two groups cross path, they then merge in a single cluster in which all walkers will follow the same path. Indeed, when a group $u$ cross path with a smaller group $v$, we have $x_{i}(t+1)=x_{j}(t)$ and $x_{j}(t+1)=x_{i}(t)$ for all $i\in g_{u}(t)$ and all $j\in g_{v}(t)$. The net flux is then $\Delta J_{u\rightarrow v}(t)=n_{u}(t)-n_{v}(t)>0$ for all walkers, which drive them all in the same direction: the smaller group goes back towards $v$ (where it came from) while the trajectory of the largest group is left unchanged.
However, this rule alone is not perfectly effective at avoiding splitting. The walkers in the first group in a cluster will not always all choose the same node at an intersection, as illustrated in the two bottom pictures of Figure 6, as the different possible nodes all have a flux equal to 0.
In this context, the Attraction rule allows walkers that split in the least chosen direction at an intersection to go backward and avoid loosing sight of the cluster.
@@
Figure 6 also shows why the baseline is not sufficient: it does not imply that walkers move forward at each step. One may try to improve this with a collective propulsion (a collective decision with the Propulsion rule instead of the Random one), but such a tactic would not ensure that groups merge when they cross each-other. This is where Alignment rule beats other approaches: walkers merge into a cluster not only if they arrive at the same node, but also if they are neighbors.
4.5 Group geography
In order to gain more insight on group formation, we display in Figure 7 the Paris street network for walkers following the Alignment strict tactic. We color nodes according to the group in which a walker that started there ends at the last step of the run. More precisely, we associate a color to each of the ten largest obtained groups. Then, for each walker that ends up in one of these groups, we color its initial location with the group color. The figure presents a typical output observed over many runs.
Figure 7 clearly shows that the walkers of a group come from a specific and limited region of the network. However these regions are rather wide, and a group does not necessarily gather all walkers that are initially located in the region.
Interestingly, even if two walkers are initially very close, or even neighbors, they often do not end in the same group. As a consequence, the regions defined by groups are entangled, like in the east of Paris (Figure 7). This is an advantage in our context: it is hard for adversary forces to predict which group a protester will join from its initial position. However, further research is needed to understand what network features lead to such behavior.
Figure 7: Discretized Paris street network with walkers following the Alignment strict tactic. We color in red the initial position of walkers that end up in the largest group at the end of the run. We do the same, with different colors, for the $10$ largest groups at the end of a run.
5 Robustness
It is crucial for protesters to form groups that resist adversary forces that may break them up. In order to explore this robustness, while sticking to our keep it simple approach, we do not consider complex models of adversary forces. Instead, we model break ups and the resulting confusion as walkers suddenly following the Random strict tactic for one step. In this way, a group located at a given node splits into smaller groups that move to neighbor nodes, in a way similar to a group of protesters targeted by adversary forces.
More formally, we perform the following experiment: we run the Alignment strict tactic for $100$ steps, then we run the Random strict tactic for one step, and we finally run the Alignment strict tactic again. We perform the same experiment with the best tactic (identified in previous sections) instead of the strict Alignment one.
Figure 8 displays the average observed scores over ten runs of each experiment, together with results of the same experiments without the Random strict tactic step. Since the coverage score is trivially increased by these experiments, which brings no useful insight, we do not display this metric.
Figure 8: Plots of robustness experiments, similar to Figure 5. We display the sprawling and gathering scores for the Alignment strict tactic and the best tactic, as well as for the experiments where all walkers perform a random move at time $100$.
The break up simulation clearly appears on both plots. The gathering score plots show that large groups reappear shortly after break ups. The two tactics do not have the same behaviour, though. The strict Alignment tactic needs some time to reach gathering scores close to its pre-break up ones (green curve in Figure 8). Instead, the best tactic is strikingly efficient in quickly recovering from break ups. Indeed, large groups do not only reappear shortly after the break up, they do so even faster than before. This brings its gathering score (black curve in Figure 8) to quickly catch up with the case with no break up (purple curve).
We also observe that, before going back to normal, the sprawling score experiences a sharp increase immediately followed by a sharp decrease. The increase is due to the fact that groups in each cluster first split into several neighbor groups, thus forming a larger cluster. Then, the new groups tend to go forward, due to the Alignment rule, and so they move away from each other, forming distinct clusters. Then, like for the gathering score, the best tactic recovers strikingly well, and faster than the Alignment strict tactic.
@@ 6 Related work
Our work is not concerned with the modeling of actual pedestrian behaviours and trajectory planning [5, 22, 20]. Instead, our goal is to design simple tactics that lead to targeted features, like flocking.
@@
Likewise, our work differs from protest models based on thresholds [17] or agent-based [16, 21, 25, 1] approaches. Indeed, these works focus on how people decide to participate in a protest; they do not deal with protester mobility.
Works in [24, 8] are much closer to ours. Not only their protesters are mobile, but authors also study adversary forces. However, they focus on the strategy these forces should follow to control the crowd, whereas we are more interested in protesters tactics.
Finally, a number of works deal with flocking behaviors of walkers in a variety of situations. We summarize their main features in Table 1 and show their limits from our perspective. We discuss them in more details below.
Table 1: Important models of collective walks found in literature that provide solutions to problems similar to ours. The $\bullet$ specifies if they succeed at solving one of the issues we deal with.
Models Network Lattice Flocking Fast
Random walk [27] $\bullet$ $\bullet$
UTS/UXS [2, 23] $\bullet$ $\bullet$
EAW [36] $\bullet$ $\bullet$ $\bullet$
SIP [18] $\bullet$
MIPS [9] $\bullet$ $\bullet$
Boids [33] $\bullet$ $\bullet$
Cucker-Smale [12] $\bullet$ $\bullet$
Robot swarm [15, 10] $\bullet$ $\bullet$ $\bullet$
SVM [34, 32] $\bullet$ $\bullet$ $\bullet$
Our work $\bullet$ $\bullet$ $\bullet$ $\bullet$
6.1 Walks on networks
First notice that our baseline is known as a coalescent random walk [11] and guarantees a non-deterministic gathering. In the field of distributed computing, researchers therefore proposed deterministic solutions for such gathering problems. Walkers then follow a common distributed algorithm to meet on any connected graph [14, 31, 3]. The main algorithms are UTS and UXS in Table 1 [2, 23].
In such works, labeling nodes during walks and knowledge of labels is in general crucial. Moreover the focus is on guarantees regarding long-term gathering. Instead, fast gathering is crucial in our context, as well as very limited memory and computation capacities.
Other works in this line focus on network exploration [35], and conclude that backtracking is to avoid, which we indeed observe with the high coverage score of the Propulsion rule.
6.2 Flocking in free space
Many papers explore flocking since the seminal Boids model [33]. These works typically consider bird swarms or pedestrian crowds that move in free space, like the continuous two-dimensional plane. Then, flocking means that a group of agents spontaneously move in the same direction [12, 29].
The work of Vicsek [34] (SVM in Table 1) notably established that alignment and noise are sufficient to produce flocking in the plan. The combination of alignment and attraction is also an effective way to gather and flock [19], as attraction avoids sprawling.
Robot swarms often target objectives very similar to ours [15]. However, they generally use the local communication model that exceeds our walkers capabilities.
Our work differs from all these works by the fact that walkers move in a network, not in free space. Therefore, they do not have a readily usable notion of direction, except in very specific cases where the network is a lattice, for instance. This makes a big difference, and some of our moving criteria may be seen as extensions of free-space concepts to networks.
6.3 Flocking and networks
Papers on flocking and networks rarely study walks in networks. They typically focus on proximity graphs of walkers: the graph of which walkers (in free space) are in sight of each other [30]. Then, the connectivity of these graphs allows to evaluate if walkers achieve flocking.
Only a few papers are concerned with flocking in networks, but then the considered networks are in general very specific. The motivation typically is to use lattices to approximate free space and study active particles [9]. For instance, the authors of [32] explore rules similar to ours in order to study their combinations on a line. They get results similar to the ones we presented in the single street case.
Finally, some algorithms solve problems close to ours on a network. For instance, [13] studies the problem of scattering walkers over a graph, when they are initially all gathered at a single node. Edge ant walk (EAW in Table 1[36] is a bio-inspired model where walkers communicate by leaving a trace of their passage, useable by other walkers. Such works require capacities that exceed the ones of our walkers.
7 Conclusion
This paper proposes a framework for exploring which basic rules are crucial for walkers to gather and flock in street networks. Our findings indicate that the Alignment strict tactic is sufficient for flocking. Combining Alignment with other tactics is even more effective, and obtained groups are very robust to break ups.
This raises many questions on group formation and dynamics over time, and on the role of street network properties in gathering and flocking. For instance, it is unclear why a group emerges at a specific location and gathers walkers from a specific region. Figure 7 is a first step in answering such questions, and it shows they are non-trivial.
Our framework may also be extended in several ways. For instance, the city modelling may be improved with more urban data, like street width or point-of-interest (POI) data (restaurants, stores, or theaters, e.g.). Such information may be used by walkers, and they open the way to studies of the impact of groups on network connectivity.
Reproducibility. We publicly provide all code and data used in this paper, our models in C and a Python code to analyse outputs: https://github.com/K-avi/protesting_on_graphs. We describe how to use it, including how to generate the obtained dataset, on the companion site: https://k-avi.github.io/protesting_on_graphs.
Acknowledgements
We warmly thank Ivan Mulot-Radojcic for his help with implementation and documentation and Nicolas Maudet for his highly valuable feedbacks. This work is funded in part by the MITI CNRS interdisciplinary program.